Three resistors, R1 = 3.00 ohm, R2 = 4.00 ohm, R3 = 5.00 ohm, are connected in series. With what single equivalent resistance Req could the series combination be replaced? If the series combination is connected to a 12.0 V battery, what is the voltage across each resistor?
Answer #1

For resistors in series, we can simply add the resistances

Req = R1 + R2 + R3

Req = 3 + 4 + 5

Req = 12 ohms

---------------------------------------------------

if this is connected to 12 V, then equivalent current will be

I = V / Req

I = 12 / 12

I = 1 Amp

this current will be same through all resistors as they are in series

so,

voltage across R1 = 1 * 3 = 3 Volts

voltage across R2 = 1 * 4 = 4 Volts

voltage across R3 = 1 * 5 = 5 Volts